A rectangular doorway in a hotel is 1.5m wide and 2.5m high. Two waiters will bring a circular table top with 2.7m in diameter. Do you think the waiters will be able to bring the table top?


1
Expert's answer
2021-05-26T03:52:57-0400

Yes! The longest possible side of the door (hypotenuse) is bigger than the diameter of the table top.

From the formula, a² + b² = c², assuming that our the width to be the a, and the height to be b, we can then calculate our hypotenuse from the above formula.

(1.5)² +(2.5)² = c²

2.25 + 6.25 =c²

Therefore, c² =8.50

C is therefore the square root of 8.5

=2.92M

The longest point (hypotenuse) 2.92M is bigger than the diameter 2.7M of the round table. Therefore, the waiters can easily slant the table through.